OpenStudy (dqswag):

The school is having a bake sale and the math club is bringing cupcakes. They have 200 total cupcakes, both chocolate and strawberry. The number of chocolate cupcakes is twenty more than twice the number of strawberry cupcakes. Which system of equations can be used to find the number of chocolate cupcakes, c, and the number of strawberry cupcakes, s?

OpenStudy (dqswag):

c + s = 200 s = 2c - 20 c - s = 200 s = 2c + 20 c + s = 200 c = 2s + 20 c = 2s c - 200 = s

OpenStudy (dqswag):

@Alisha007

OpenStudy (anonymous):

Not 100% positive on this one, so if anyone finds a mistake let me know but if im not mistaken it should be c + s = 200 s = 2c - 20

OpenStudy (dqswag):

It was c + s = 200 c = 2s + 20

OpenStudy (dqswag):

But good try